# Ensemble-Based Data Assimilation for Material Model Characterization in High-Velocity Impact

### v1

@@ 5.5 EnKF Inversion Results
Case 4: strongly biased initial guess, $\bm{u}_{0}=[0.0325,1.1845,3.85]^{\text{T}}$, $N_{\rm{o}}=20$
Prior $3.24\times 10^{-2}\pm 1.01\times 10^{-3}$ +149.50+149.50 $1.18\pm 4.64\times 10^{-2}$ +149.89+149.89 $3.84\pm 1.58\times 10^{-1}$ +149.38+149.38
Posterior $2.05\times 10^{-2}\pm 4.71\times 10^{-5}$ +57.63+57.63 $1.07\pm 4.28\times 10^{-2}$ +126.43+126.43 $1.79\pm 2.15\times 10^{-2}$ +16.50+16.50
@@ 6 Discussion
Understanding and predicting the performance of materials under HVI play a critical role in both fundamental research and practical applications. Experimental testing, however, is often limited due to cost, time, and technical challenges. High-fidelity simulations provide a valuable alternative, but require accurate material models and reliable model parameters, including those for plasticity, fracture, and EOS. Traditionally, these parameters are estimated by manually tuning them until a visually acceptable agreement is achieved with the selected experimental results. For example, plasticity and fracture parameters can be obtained from Split-Hopkinson pressure and tension bar experiments, while EOS parameters are typically determined from plate impact tests. Nevertheless, such experiments are labor- and time-intensive, and the manual tuning process can introduce additional uncertainties. In this work, we propose an alternative approach based on EnKF: directly assimilating HVI observations into SPH simulations to characterize material properties. This strategy bypasses the need for separate calibration experiments and manual parameter tuning and enhances the predictive capability of simulations.
@@
In addition, a strongly biased initial guess can delay or even inhibit convergence, including parameters that are demonstrably sensitive to the data. In Case $4$, as shown in Fig. 9, both $C$ and $\gamma_{0}$ initially converge toward the truth during the early iterations; however, they subsequently plateau, showing a consistent positive bias. Conversely, $D_{4}$ remains weakly identifiable and may continue to diverge. The ”drift-then-stall” phenomenon manifests when the initial ensemble is positioned significantly distant from the true solution’s basin of attraction. The unrepresentative sample covariance results in suboptimal Kalman gains and, in conjunction with premature variance reduction, constrains updates along skewed directions. The overall outcome is a significant decrease in posterior dispersion without corresponding improvements in accuracy, indicative of limited observability and prior misspecification rather than instability of the filter. Practically, judicious initialization and prior design, together with adaptive inflation or periodic resampling, and, when possible, additional observations, can mitigate this sensitivity. Overall, this stress test underscores the robustness of the EnKF to severe prior misspecification: stability is preserved, prediction error decreases, and the method highlights which parameters require improved observability or prior mass to achieve full recovery.
@@ 7 Concluding Remarks
In this study, we have developed an EnKF-based inversion framework to calibrate material models used in HVI simulations. The framework has been applied to identify the JC plasticity, fracture, and Grüneisen EOS parameters in SPH simulations of a steel ball impacting an AZ31B magnesium plate, using time-series measurements of back-face deflection. A parameter screening analysis has been first performed to reduce the dimensionality of the unknown input space. Subsequently, four cases have been designed to examine the influence of initial guesses and the quantity of observational data on the calibration outcomes. From these investigations, the following conclusions can be drawn:
@@
  2. 2.
    Using the same amount of data, the EnKF framework fails to identify the insensitive parameter $D_{4}$ in the JC fracture model. Its estimates retain large errors (greater than $10\%$) and a wide ensemble spread comparable to that of the initial ensemble, consistent with its low sensitivity.
@@
    An accurate initial ensemble mean and sufficient observational data are critical for EnKF inversion. When there is an increase in the initial ensemble bias or a reduction in the quantity of observational data, the efficiency and accuracy of EnKF estimation diminish.
These conclusions highlight three critical implications: first, sensitivity screening is essential for differentiating between parameters that are identifiable and those that are not; second, adaptive inflation serves as a significant protective measure against inaccurate initial assumptions; and third, to attain robust and uncertainty-conscious calibration from a solitary high-velocity impact experiment, it is imperative to maintain a minimum threshold of observational information content, which can be accomplished either by augmenting sensor density or by integrating complementary diagnostic methods.
@@ 8 Acknowledgement
R.J. and X.S. gratefully acknowledge the support of the National Science Foundation (NSF), USA under Grant No. $2429424$, and the support of the University of Kentucky, USA through the Faculty Start-up Fund. G.W. gratefully acknowledges the support of the National Natural Science Foundation of China (52301336), Science and Technology Development Fund of Macau S.A.R. (0048/2025/ITP1), and University of Macau (SRG2025-00004-FST). R.J. and X.S. appreciate the insightful discussions with Dr. K. T. Ramesh, Dr. Suhas Eswarappa Prameela, Dr. Pinkesh Malhotra, and Dr. Justin Moreno.

### v2

@@ 6.5 EnKF Inversion Results
Case 4: strongly biased initial guess, $\bm{u}_{0}=2.5\,\bm{u}_{\text{true}}$, $N_{\rm{o}}=20$
Prior $3.24\times 10^{-2}\pm 1.26\times 10^{-3}$ +149.5 $1.18\pm 9.10\times 10^{-2}$ +149.89 $3.84\pm 3.25\times 10^{-1}$ +149.38
Posterior $1.31\times 10^{-2}\pm 2.00\times 10^{-4}$ +0.81+0.81 $1.12\pm 8.28\times 10^{-2}$ +136.94+136.94 $1.51\pm 4.46\times 10^{-2}$ 2.16-2.16
Figure 12: Comparison of back-face deflection profiles at $t=10.0,11.0$, and $12.0~\mu s$: (a) Case 1, (b) Case 2, (c) Case 3, and (b) Case 4.
To visually assess the quality of the calibration, Fig. 12 compares the back-face deflection profiles at the three representative time instants for Cases 1-4. Each figure overlays the baseline deflection computed at the true parameter values, the prediction from the initial guess (Prior), and the prediction from the calibrated model (EnKF Posterior). In Cases 2 and 4, where the prior prediction (dashed lines) shows a clear deviation from the baseline (solid lines), the calibrated profiles (dotted lines) closely align with the ground truth, demonstrating the effectiveness of the filter. In Cases 1 and 3, where the initial guess visually appears similar to the truth, the EnKF framework nevertheless accurately identifies the sensitive parameters $C$ and $\gamma_{0}$, providing calibrated results suitable for subsequent HVI simulations.
[…]
Figure 13: Comparison of the post-calibration fitting error against the empirical sensitivity bound at: (a) $t=10.0~\upmu s$, (b) $t=11.0~\upmu s$, and (c) $t=12.0~\upmu s$.
@@ 7 Discussion
Understanding and predicting the performance of materials under HVI play a critical role in both fundamental research and practical applications. However, experimental testing is often limited due to cost, time, and technical challenges. High-fidelity simulations provide a valuable alternative, but require accurate material models and reliable model parameters, including those for plasticity, fracture, and EOS. Traditionally, these parameters are estimated by manually tuning them until a visually acceptable agreement is achieved with the selected experimental results. For example, plasticity and fracture parameters can be obtained from Split-Hopkinson pressure and tension bar experiments, while EOS parameters are typically determined from plate impact tests. Nevertheless, such experiments are labor- and time-intensive, and the manual tuning process can introduce additional uncertainties. In this work, we propose an alternative approach based on EnKF: directly assimilating HVI observations into SPH simulations to characterize material properties. This strategy bypasses the need for separate calibration experiments and manual parameter tuning and enhances the predictive capability of simulations.
A key consideration for SPH-based inverse modeling is computational cost. Using a simple impact-inspired example, we have verified that EnKF is at least one order of magnitude more efficient than a traditional MCMC method in terms of the number of forward evaluations. This advantage can be further amplified by the inherent parallelism of EnKF. Within each EnKF iteration, the ensemble forward simulations are independent and can be executed concurrently, which is particularly beneficial for computationally expensive HVI simulations. In our numerical HVI studies, each SPH run takes approximately 10 minutes. If EnKF converges within $10$ iterations with an ensemble size of $100$, this corresponds to roughly 1,0001{,}000 forward runs in total. With $100$ parallel cores, this translates to about $100$ minutes of wall-clock time, or approximately $10$ hours using $16$ cores. In contrast, an RWM-MCMC implementation requiring $\mathcal{O}(10^{4})$ sequential steps (without surrogate models) would take on the order of $69$ days. These projections underscore that standard MCMC methods remain computationally prohibitive for this class of problems unless surrogate models are introduced, whereas EnKF offers a practical pathway for calibrating expensive black-box models by leveraging parallel computing to reduce wall-clock time.
@@
To further evaluate the scalability of the EnKF framework, we consider a higher-dimensional HVI problem parameterized by sensitive unknowns $\bm{u}=[A,C,D_{1},D_{2},C_{s},\gamma_{0}]^{\text{T}}$. All other simulation settings are identical to those in Case 1, and detailed calibration results are reported in Appendix B. The main finding is that $A$, $C$, $C_{s}$, and $\gamma_{0}$ converge to their true values, although they require more iterations than in Case 1. In contrast, $D_{1}$ and $D_{2}$ do not converge to their true values and retain relatively large ensemble spreads. Examination of the EnKF performance in the observation space suggests that this non-convergence and large uncertainty for $D_{1}$ and $D_{2}$ may arise from non-uniqueness with respect to the chosen observables: different combinations of these fracture parameters yield nearly indistinguishable back-face deflection profiles. Multiple realizations of the initial ensemble are tested, and this behavior persists consistently. This phenomenon underscores the ill-posed nature of HVI inverse problems in higher-dimensional parameter spaces, where distinct fracture mechanisms can compensate for one another to produce similar global responses.
This six-dimensional problem stands in sharp contrast to the three-parameter cases in the main text, where sensitive material parameters consistently converge to the true values with small uncertainties regardless of the initial prior bias. We emphasize again that OAT is used here only as a local sensitivity metric for pre-screening. Low RMSD values from the OAT analysis indicate low local sensitivity and non-uniqueness of the corresponding parameters. However, high RMSD values alone cannot guarantee uniqueness or identifiability. Therefore, in practical applications where the ground truth is unknown, a multi-start strategy—performing calibration from several distinct initial guesses—is essential to verify the robustness of convergence and to diagnose potential unidentifiability through persistently large ensemble variance.
@@ 8 Concluding Remarks
In this study, we have developed an EnKF-based inversion framework to calibrate material models used in HVI simulations. The framework has been applied to identify representative JC plasticity, fracture, and Grüneisen EOS parameters in SPH simulations of a steel ball impacting an AZ31B magnesium plate, using time-series measurements of back-face deflection. While other parameters are fixed based on literature or conventional tests following a sensitivity screening, the selected subset are used to demonstrate the framework’s ability to recover sensitive parameters and diagnose non-identifiable ones. Subsequently, four three-dimensional and one six-dimensional cases have been designed to examine the influence of initial guesses, the quantity of observational data, and the number of unknowns on the calibration outcomes. From these investigations, the following conclusions can be drawn:
@@
    Compared to traditional derivative-free Bayesian MCMC methods, the EnKF framework is at least one order of magnitude more efficient in terms of forward model evaluations. By leveraging parallel computing, the wall-clock time for calibration is reduced from months to hours, making it a practically feasible pathway for high-fidelity, black-box HVI simulations.
  2. 2.
@@
    The framework demonstrates robustness against extreme prior misspecification through a parameter rejuvenation strategy. Under $150\%$ initial bias, this strategy prevents premature filter collapse, allowing sensitive parameters to migrate from outside the initial ensemble spread toward the true values. Without such intervention, the filter converges to incorrect estimates with small uncertainties.
  4. 4.
    The high-dimensional inversion study highlights the fundamental challenge of equifinality in HVI calibration: multiple parameter combinations can produce virtually indistinguishable responses. It shows that, while the proposed framework delivers reliable predictive performance, parameter-level uniqueness in higher-dimensional settings can be compromised by strong correlations among influential parameters. Such non-uniqueness is reflected by persistently large ensemble uncertainties across EnKF iterations, and it may be alleviated by incorporating additional and more informative observations.
@@ 9 Acknowledgement
R.J. and X.S. gratefully acknowledge the support of the National Science Foundation (NSF), USA under Grant No. $2429424$, and the support of the University of Kentucky, USA through the Faculty Start-up Fund. G.W. gratefully acknowledges the support of the National Natural Science Foundation of China (52301336), Science and Technology Development Fund of Macau S.A.R. (0048/2025/ITP1), and University of Macau (SRG2025-00004-FST). R.J. and X.S. appreciate the insightful discussions with Dr. K. T. Ramesh, Dr. Suhas Eswarappa Prameela, Dr. Pinkesh Malhotra, and Dr. Justin Moreno. The suggestions of the anonymous reviewers have helped to improve the quality and scope of this work.
Appendix
Appendix A Baseline EnKF Performance under Extreme Bias
To demonstrate the necessity of the rejuvenation strategy introduced in Section 4.2, we present here the results of Case 4 using the EnKF implementation only with RTPS inflation under the same extreme initialization ($\bm{u}_{0}\approx 2.5\,\bm{u}_{\text{true}}$). As illustrated in Fig. 15, the standard filter exhibits a characteristic “drift-then-stall” behavior. The ensemble variance collapses rapidly within the first 3-5 iterations (indicated by the narrowing grey band), causing the filter to lock into an incorrect solution before the mean can migrate to the true values. Quantitatively, while the initial discrepancies were approximately $150\%$ for all parameters, the standard EnKF could only reduce the errors of the sensitive parameters $C$ and $\gamma_{0}$ to $57.63\%$ and $16.50\%$, respectively. This contrasts starkly with the results of Case 4 using the triggered rejuvenation method, which achieved errors of $0.81\%$ and $2.16\%$. Furthermore, the standard filter reports a deceptively small posterior standard deviation for these biased estimates, a phenomenon known as “false confidence”. This comparison highlights that without parameter rejuvenation, the filter is unable to recover from strong prior biases even when parameters are identifiable.
Figure 15: EnKF results for Case 4 without parameter rejuvenation: (a)-(c) Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d) Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.
Appendix B EnKF Inversion For a High-Dimensional Case
We introduce a high-dimensional problem: an expanded parameter vector defined as $\bm{u}=[A,C,\allowbreak D_{1},D_{2},\allowbreak C_{s},\gamma_{0}]^{\text{T}}$. In this scenario, the calibration scope is expanded to six parameters, selecting the most sensitive ones from each material model: $A$ and $C$ (plasticity), $D_{1}$ and $D_{2}$ (fracture), and $C_{s}$ and $\gamma_{0}$ (EOS). The initial ensemble is generated with a $25\%$ bias (${\bm{u}}_{0}=0.75\bm{u}_{\rm true}$), consistent with the setup in Case 1, and the same observation set ($N_{o}=20$) is utilized. The inversion results for this case are presented in Fig. 16 and summarized in Table 7.
The parameters $A$, $C$, $C_{s}$, and $\gamma_{0}$ converge toward their true values, albeit requiring more iterations than in Case 1 as expected. Table 7 shows that the relative errors for $A$ and $C_{s}$ are reduced to $0.07\%$ and $0.23\%$, respectively. $C$ and $\gamma_{0}$ achieve final errors of approximately $1\%$. The posterior uncertainties for these four parameters are also substantially reduced, indicating strong identifiability. In contrast, the fracture parameters $D_{1}$ and $D_{2}$ do not converge to their true values. Although they were identified as sensitive in the OAT analysis in Section 6.4, they collapse to incorrect values and retain relatively large ensemble spreads, with residual biases of $21.6\%$ and $21.9\%$, respectively. This outcome highlights that OAT, as a local sensitivity metric, does not guarantee identifiability in a higher-dimensional inverse problem.
To understand why $D_{1}$ and $D_{2}$ do not converge to their true values, we examine the performance of the EnKF framework in the observation space, as shown in Fig. 17. Despite clear deviations in parameter space, the calibrated model achieves excellent predictive accuracy in the observable. Specifically, Fig. 17(a) shows that the RMSE of the observation vector decreases rapidly and stabilizes at a negligible level ($\sim 10^{-4}$ mm) after about ten iterations. Fig. 17(b) further confirms that the time-resolved error in the back-face deflection remains consistently small throughout the impact process. The posterior back-face deflection profiles in Fig. 17(c) are in excellent agreement with the true response, effectively correcting the initial bias at the output level.
Taken together, Figs. 16 and 17 may provide an evidence of non-uniqueness with respect to the chosen observable. They show that different parameter sets—namely, the true set versus the calibrated set with biased $D_{1}$ and $D_{2}$—can produce virtually indistinguishable back-face deflection profiles. This indicates that, in the present regime, the available back-face deflection data, while sufficient to constrain stiffness and compressibility (dominated by $A,C,C_{s},\gamma_{0}$), remain compatible with multiple combinations of the fracture initiation and evolution parameters ($D_{1},D_{2}$). This finding underscores an important limitation: although the EnKF framework reliably ensures predictive accuracy for the observable quantity, the physical interpretability of individual parameters hinges on identifiability, and strict uniqueness may require additional diagnostic data (e.g., perforation hole diameter, residual velocity, or back-face at more time instances) to break such parameter correlations.
Figure 16: EnKF inversion results for the high-dimensional case with 6 unknown parameters. (a)-(f) Iteration histories of the ensemble mean (blue line) and spread (gray area) for parameters $A,C,D_{1},D_{2},C_{s},\gamma_{0}$. The red dashed line indicates the true value.
Figure 17: Assessment of predictive capability for the high-dimensional case in the observation space: (a) Iteration history of the observation RMSE. (b) Time-resolved average error of the back-face deflection before ($l=0$) and after ($l=20$) assimilation. (c) Comparison of back-face deflection profiles at $t=10.0,11.0$, and $12.0~\mu s$.
Table 7: Summary of EnKF results for the high-dimensional case.
Parameter Truth Prior mean $\pm$ std Posterior mean $\pm$ std Error [%]
AA [MPa] $2.25\times 10^{2}$ $1.68\times 10^{2}\pm 1.32\times 10^{1}$ $2.25\times 10^{2}\pm 3.58$ +0.07+0.07
$C$ $1.30\times 10^{-2}$ $9.74\times 10^{-3}\pm 9.55\times 10^{-4}$ $1.29\times 10^{-2}\pm 4.98\times 10^{-4}$ 1.06-1.06
$D_{1}$ $3.50\times 10^{-1}$ $2.64\times 10^{-1}\pm 2.69\times 10^{-2}$ $2.75\times 10^{-1}\pm 1.17\times 10^{-2}$ 21.60-21.60
$D_{2}$ $6.03\times 10^{-1}$ $4.57\times 10^{-1}\pm 4.58\times 10^{-2}$ $4.71\times 10^{-1}\pm 1.91\times 10^{-2}$ 21.90-21.90
$C_{s}$ [m/s] $4.520\times 10^{3}$ $3.38\times 10^{3}\pm 3.37\times 10^{2}$ $4.51\times 10^{3}\pm 1.53\times 10^{1}$ 0.23-0.23
$\gamma_{0}$ $1.54$ $1.14\pm 1.21\times 10^{-1}$ $1.52\pm 4.03\times 10^{-2}$ 1.01-1.01